# Algorithms for Markov Binomial Chains

### v1

@@ 6.1. A second algorithm, now without Gauss
Before proving the result, we observe that 5 gives us a characterization of the nonzero terms in $\boldsymbol{P}$. In this way, for SIR binomial chains, the expression from 7 is reduced to what is shown in the next lemma.
@@
Hence, due to our choice of order on the states, together with 5, we get that $\boldsymbol{I}-\boldsymbol{Q}$ is upper triangular. This enables the computation of $\boldsymbol{k^{A}}$ by means of back substitution, which requires only a quadratic number of arithmetic operations.
@@
It remains to argue that $\boldsymbol{P}$ can be precomputed in $O(N^{4})$. Unfortunately, computing every entry of $\boldsymbol{P}$ using Equation 15 seems to require time $O(N)$. Indeed, while exponentiation can be realized using a logarithmic number of operations via iterated squaring, we are not aware of an algorithm to compute factorials using a sublinear number of multiplications. [^1] We do know that, in the Turing-machine model of computation, one can obtain seemingly better bounds [Bor85]. However, those depend on the complexity of multiplying numbers being a function of the amount of bits used to encode them.
@@ 6.2. Precomputation of the transition probabilities using dynamic programming
Proof 6.1.
Note that $P(\boldsymbol{m},\boldsymbol{n}),P(m_{1}-1,m_{2},m_{3}+1,\boldsymbol{n})>0$ by 5 and our assumptions. We first consider the binomial coefficients of (29) and (30).
@@
We observe that 6.1 clearly terminates because all for-loops are bounded and there are no jump statements in the code. Regarding its time complexity, we note that we only ever nest at most $4$ for-loops and $P(\boldsymbol{m},\boldsymbol{n})$ is never used (i.e. computed naively in $O(N)$ steps) within a for-loop nesting of depth $4$.
@@
The worst-case time complexity of 6.1 is $O(N^{4})$ in the Blum-Shub-Smale model.
@@
Let $X$ be as computed by 6.1. Then, $X(\boldsymbol{m},\boldsymbol{n})=P(\boldsymbol{m},\boldsymbol{n})$ for all pairs of states $\boldsymbol{m},\boldsymbol{n}$ such that $n_{1}\leq m_{1}$ and $m_{3}\leq n_{3}\leq m_{2}+m_{3}$.
Proof 6.2.
@@ 7.1. Stochastic counter machines
A stochastic counter machine (or SCM) in dimension $k$ is a tuple $\mathcal{M}=(Q,q_{0},\boldsymbol{c_{0}},T)$ where $Q$ is a finite set of (control) states, $q_{0}\in Q$ is the initial state, $\boldsymbol{c_{0}}\in\mathbb{N}^{k}$ is the initial value of the counters (to which we sometimes refer using Greek letters), and $T$ is a finite set of tuples $(q,g,u,q^{\prime},p)$ where: $q,q^{\prime}\in Q$, the guard $g$ is a matrix-vector pair $\boldsymbol{A}\in\mathbb{Q}^{n\times d},\boldsymbol{b}\in\mathbb{Q}^{n}$, the udapte $u$ is also a matrix-vector pair $\boldsymbol{U}\in\mathbb{Q}^{d\times d},\boldsymbol{r}\in\mathbb{Q}^{d}$, and $p\in\mathbb{Q}_{\geq 0}$.
@@ 9. Conclusion and future work
We have started the study of binomial chains through the lens of formal methods. In this work, we provided two main theoretical results. First, we established that acyclic binomial chains almost surely terminate. Second, we proved the the problem of approximating the time to termination is in PSPACE and gave a direct algorithm for the exact problem (ignoring the complexity of arithmetic and the annoyances of irrational numbers). We also provided a more pragmatic approach in the form of an encoding into the PRISM language. For this last approach we also presented some experiments. Based on the empirical results, we can conclude that state-of-the-art probabilistic model checkers are not (yet) powerful enough to deal with epidemiological models like the one proposed in [AWS+21].Indeed, while Modest is capable of handling simple probabilistic queries for realistic populations, it still seems to struggle with quantitative queries such as the expected end of epidemic property. In this direction, more research is needed to find, for instance, good abstractions.

### v2

@@ 7.1. A second algorithm, now without Gauss
Before proving the result, we observe that 2 gives us a characterization of the nonzero terms in $\boldsymbol{P}$. In this way, for SIR binomial chains, the expression from 7 is reduced to what is shown in the next lemma.
@@
Hence, due to our choice of order on the states, together with 2, we get that $\boldsymbol{I}-\boldsymbol{Q}$ is upper triangular. This enables the computation of $\boldsymbol{k^{A}}$ by means of back substitution, which requires only a quadratic number of arithmetic operations.
@@
It remains to argue that $\boldsymbol{P}$ can be precomputed in $O(N^{4})$. Unfortunately, computing one single entry of $\boldsymbol{P}$ using Equation 12 seems to require time $O(N)$, meaning that a naive enumeration of pairs of states $(m_{1},m_{3})$ and $(n_{1},n_{3})$ with $P((m_{1},N-m_{1}-m_{3}),(n_{1},N-n_{1}-n_{3},n_{3}))$ computed for each of them results in a total of $O(N^{5})$. Indeed, while exponentiation can be realized using a logarithmic number of operations via iterated squaring, we are not aware of an algorithm to compute factorials using a sublinear number of multiplications. [^1] We do know that, in the Turing-machine model of computation, one can obtain seemingly better bounds [Bor85]. However, those depend on the complexity of multiplying numbers being a function of the amount of bits used to encode them.
@@ 7.2. Precomputation of the transition probabilities using dynamic programming
Proof 7.1.
Note that $P(\boldsymbol{m},\boldsymbol{n}),P(m_{1}-1,m_{2},m_{3}+1,\boldsymbol{n})>0$ by 2 and our assumptions. We first consider the binomial coefficients of (31) and (32).
@@
We observe that 7.1 clearly terminates because all for-loops are bounded and there are no jump statements in the code. Regarding its time complexity, we note that we only ever nest at most $4$ for-loops and $P(\boldsymbol{m},\boldsymbol{n})$ is never used (i.e. computed naively in $O(N)$ steps) within a for-loop nesting of depth $4$.
@@
The worst-case time complexity of 7.1 is $O(N^{4})$ in the Blum-Shub-Smale model.
@@
Let $X$ be as computed by 7.1. Then, $X(\boldsymbol{m},\boldsymbol{n})=P(\boldsymbol{m},\boldsymbol{n})$ for all pairs of states $\boldsymbol{m},\boldsymbol{n}$ such that $n_{1}\leq m_{1}$ and $m_{3}\leq n_{3}\leq m_{2}+m_{3}$.
Proof 7.2.
@@ 8.1. Stochastic counter machines
A stochastic counter machine (or SCM) in dimension $k$ is a tuple $\mathcal{M}=(Q,q_{0},\boldsymbol{c_{0}},T)$ where $Q$ is a finite set of (control) states, $q_{0}\in Q$ is the initial state, $\boldsymbol{c_{0}}\in\mathbb{N}^{k}$ is the initial value of the counters (to which we sometimes refer using Greek letters), and $T$ is a finite set of tuples $(q,g,u,q^{\prime},p)$ where: $q,q^{\prime}\in Q$, the guard $g$ is a matrix-vector pair $\boldsymbol{A}\in\mathbb{Q}^{n\times d},\boldsymbol{b}\in\mathbb{Q}^{n}$, the update $u$ is also a matrix-vector pair $\boldsymbol{U}\in\mathbb{Q}^{d\times d},\boldsymbol{r}\in\mathbb{Q}^{d}$, and $p\in\mathbb{Q}_{\geq 0}$.
@@ 10. Conclusion and future work
We started the study of binomial chains through the lens of formal methods. In this work, we provided two main theoretical results. First, we established that acyclic binomial chains almost surely terminate. Second, we proved the the problem of approximating the time to termination is in PSPACE and gave a direct algorithm for the exact problem (ignoring the complexity of arithmetic and the annoyances of irrational numbers). Unfortunately, extending the algorithm from section 7 to general binomial chains seems hard. The key result enabling 7.1 to compute transition probabilities in $O(N^{4})$ was 22 and it made use of the explicit formula for transition probabilites we had manually derived for SIR binomial chains. While similar results could be obtained by hand for fixed transfer matrices, it is unclear to us what a (meta)result for general binomial chains would look like.
We also provided a more pragmatic approach in the form of an encoding into the PRISM language. For this last approach we also presented some experiments. Based on the empirical results, we can conclude that state-of-the-art probabilistic model checkers are not (yet) powerful enough to deal with epidemiological models like the one proposed in [AWS+21]. Indeed, while Modest is capable of handling simple probabilistic queries for realistic populations, it still seems to struggle with quantitative queries such as the expected end of epidemic property. In this direction, more research is needed to find, for instance, good abstractions.